Chapter 2 Economic Models 25 3. Graph the functions (o) y -x2+5x-2 (b) y x+5x-2 with the set of values-5- 5 constituting the domain. It is well known that the sign of the coefficient of the x2 term determines whether the graph of a quadratic func- tion will have a hill or a valley. On the basis of the present problem, which sign is associated with the hill? Supply an intuitive explanation for this. Please write out steps and explain the graph! I’m really confused !
Answer #1

Solution:

40 -30 20 -10 20 30 40 20

The green graph shows y=x^2+5x-2 and purple graph shows y=-x^2+5x-2. The graph is constructed by putting values of x between -5 and 5 and corresponding values of y and smoothening out the curve.

As the graph shows, the quadratic equation having minus sign with x^2 will have hill and that having positive sign with x^2 will have valley.